OpenStudy (anonymous):

Is 1+cos^2(7x) equal to sin^2(7x)? Or am I off base here?

OpenStudy (debbieg):

\(\Large \sin^2\alpha + \cos^2\alpha=1\) for all real numbers. So it follows that: \(\Large \sin^2\alpha =1\color{red}- \cos^2\alpha\) But I don't see how you would get \(\Large \sin^2\alpha =1\color{red}+ \cos^2\alpha\) ??

OpenStudy (anonymous):

Ah, you are right. Trying to use u substitution for an inverse trig problem, but I think I figured it out now. Thank you.

OpenStudy (debbieg):

You're welcome. :)